OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5

OpenStudy (anonymous):

2x + 3y = 4 x - 2y = -5 ----------- multiply the bottom equation by -2 2x + 3y = 4 -2x +4y = 10 ------------- add vertically 2x-2x + (3y+4y) =4 + 10 7y = 14 y = 14/7 y = 2 plug y = 2 into the equation 2x + 3y = 4 for y and solve for x.

OpenStudy (anonymous):

thank you!:)

OpenStudy (anonymous):

you're welcome